OpenStudy (anonymous):

how to find the restriction of 9x^2-25/3x-5

OpenStudy (jdoe0001):

what do you mean by restriction?

OpenStudy (anonymous):

you need to find what values x can not assume, there are a few rules for this, amoung them are that sqaure roots can not be negative, and (for example here) the denominator can not have a 0 in a fraction.

OpenStudy (anonymous):

i simplify the problem to 3x+5 but then it asks for restrictions either 5/3 or -5/3

OpenStudy (jdoe0001):

what's a restriction?

OpenStudy (anonymous):

i think it is what x cannot equal

OpenStudy (anonymous):

but i have no idea how to find it

OpenStudy (anonymous):

yes, remember that the decominator can NOT be zero so you need to when 3x + 5 = 0 after you find that x value that is what x can not assume

OpenStudy (anonymous):

oh so it equals -5/3 then?

OpenStudy (anonymous):

do you always find the restriction after you simplify the problem?

OpenStudy (jdoe0001):

yes, what "x" mustn't be in this case you have a fraction, a rational, 0/5 = 0 0/215 = 0 5/0 = undefined 215/0 = undefined so, you cannot have a 0 in the denominator for a fraction to yield a valid value

OpenStudy (jdoe0001):

you don't apply the restriction to the simplified version, it always applies to the original one

OpenStudy (anonymous):

you find the restriction BEFORE you simply the problem

OpenStudy (anonymous):

from the numerator?

OpenStudy (anonymous):

set the numerator equal to zero?

OpenStudy (anonymous):

if you simply the problem you find that it is just the equation of a line. this type of problem has a removeable discontinuity (AKA hole). so it would look something like this. 

OpenStudy (anonymous):

the only restriction on fractions is that the denominator can not equal 0. the numerator can be 0 though.

OpenStudy (anonymous):

i see what your saying, thanks alot